OpenStudy (anonymous):

Solve and Check the following equations for the indicated variable: a. ln (3t + 8) = 2

OpenStudy (anonymous):

Take antilog on both sides... so our equation becomes 3t+8 = exp(2) therefore, t=(exp(2)-8)/3